Invest 40,000.00 for one year at 10% interest and 8% interest. $3360 for interest total. What was the amount of interest for each rate?
Answer by stanbon(75887)

Invest 40,000.00 for one year at 10% interest and 8% interest. $3360 for interest total. What was the amount of interest for each rate?
-------------------
Total;;;;;;;: x + y = 40000
Interest: 0.10x+0.08y = 3300
-----------------------------------
Multiply thru 1st by 0.1 to get:
0.10x + 0.10y = 4000
-----------
Subtract that from the "Interest" equation:
-0.02y = -700
y = $35000 (amount invested at 8%
Interest at this rate was $2800
----------------
Then x = $5000 (amount invested at 10%)
Interest at this rate was $500
============================================
Cheers,
Stan H.